# Hybrid Near-field and Far-field Localization with Holographic MIMO

### v1

@@ I Introduction
Traditional HMIMO-based localization works mainly fall into two categories: far-field (FF)[5, 6, 7] and near-field (NF) localization schemes[8, 9, 10]. In the FF region, the distances between the HMIMO and users are sufficiently large, and the signal channel is effectively described using a plane wave model. In [5], the authors considered FF user localization aided by millimeter wave (mmWave) HMIMO, where the users receive the signal transmitted by the BS to locate themselves. A maximum likelihood (ML) location estimator and two sub-optimal estimators with lower complexity were derived. HMIMO-based localization without base stations (BSs) was also investigated in [6], where the user transmitted the signal and received the signal reflected from the HMIMO. In [11], the authors considered the scenario that both the LOS path and the reflected path via the HMIMO exist simultaneously and proposed a low-complexity ML estimator.
@@
HMIMO-based NF localization schemes have attracted much attention recently due to an increasing interest in large-scale HMIMO. Specifically, to improve beam control capability, the scale of HMIMO is enlarged, leading to the expansion of the NF region[12]. Different from the FF case, the spherical wave model has to be employed to characterize the NF channel because the signal transmission characteristics of the NF are determined by both the range and angle of the user location. In [10], the authors considered the scenario where the LOS path between the BS and the user was blocked and used HMIMO to construct virtual line-of-sight (VLOS) paths between the users and the BS. A localization scheme targeting NF wireless users was investigated. In [9], a HMIMO functioning as a lens receiver was utilized to achieve NF localization. The authors in [8] developed a received signal strength (RSS)-based localization method for NF users and a RIS phase shift adjustment method to enhance accuracy.
@@
In practical scenarios, users may exist in both the NF and FF regions of the HMIMO. Existing algorithms for pure NF or FF users are therefore not always applicable. Specifically, FF localization methods suffer from low localization accuracy for NF user localization due to the significant error in describing NF channels using the plane wave model. In contrast, NF localization techniques are not ideal for FF users because the coupling of range and angle in the spherical model brings extra complexity, which has two main effects: first, it prolongs the running time of the algorithm; second, the more complex NF model will hinder convergence of the method and lead to accuracy degradation. In addition, the localization gain for the FF users brought by the NF model is limited. Hence, a localization scheme is required to adapt to the localization of both the NF and FF users.
@@
Several challenges in this scenario need to be addressed. First, it is unknown whether each user or scatter is in the NF or the FF region, which leads to difficulty in designing the localization algorithm. Second, the channels of different users are coupled due to the existence of common scatters and shared RIS phase shifts. To tackle the above challenges efficiently, a hybrid NF and FF localization algorithm is proposed, in which we first sample the NF and FF regions separately, and then compare the received signals with the reconstructed signals at the sampled locations to locate the users and scatters. The locations of the users and scatters are jointly estimated by considering the channel coupling effect. Then, a RIS phase shift optimization problem is formulated to minimize the sum of the Cramér-Rao bounds (CRBs) of the user localization errors. A complex circle manifold-based method is proposed to solve the optimization problem with the coupled CRBs.
@@
The rest of this paper is organized as follows. The localization scenario, the signal model, and the localization protocol are provided in Sec. II. In Sec. III, we propose the localization algorithm. Then, the RIS phase shift optimization algorithm is designed in Sec. IV. The performance of the proposed method is analyzed in Sec. V. In Sec. I, we present the simulation results and discussions. Finally, conclusions are drawn in Sec. VII.
@@ II-A Localization Scenario
As shown in Fig. 1, we consider a localization scenario consisting of $K$ users, a RIS, and a BS. Users and the BS are each equipped with a single antenna. The RIS containing $N=N_{1}\times N_{2}$ elements is placed on the Y-Z plane, where the center of the RIS is at the origin of the coordinate system. There also exist obstacles, which reflect the signals and create multipath channels. We assume the line-of-sight (LOS) path between each user and the BS is blocked, and there only exist reflection paths via the RIS between the users and the BS.
During the localization process, each user sends a narrowband signal to the RIS, and the RIS reflects it to the BS. To avoid multi-user interference, frequency division multiplexing (FDM) is applied to transmit signals from different users, i.e. the $k$-th user sends the signal over frequency $f_{k}$, and we have $f_{k}\neq f_{k}^{\prime}$ ($k\neq k^{\prime}$). The BS then estimates the locations of the users using the received signals. We assume the locations of the RIS and the BS are known, while the locations of the users and scattering centers in the obstacles [13] are unknown. Whether the users are located in the NF or the FF regions of the RIS is also unknown.
@@ II-B1 Channel Models for the NF Region
$\bm{b}(R_{k},\theta_{k},\phi_{k})=\left[\exp\left(-j\dfrac{2\pi}{\lambda}d_{k1}^{t}\right),...,\exp\left(-j\dfrac{2\pi}{\lambda}d_{kN}^{t}\right)\right]^{T}.$ (4)
@@
$\Delta\varphi(\bm{p})=\mathop{\max}\limits_{n}\frac{2\pi}{\lambda}\big{(}d_{n}^{t}-(R-y_{n}\sin\theta\sin\phi-z_{n}\cos\theta)\big{)},$ (5)
@@
Similarly, we model the paths via the scatters. Specifically, when the $l$-th scatter is located in the NF region, the path between the $l$-th scatter and the $n$-th RIS element is given by
@@ II-B2 Channel Models for the FF Region
where $\bm{a}(\theta_{k},\phi_{k})$ is the steering vector in the FF region and is given as[17]
@@
In (7), $\bm{D}_{FF}$ represents the FF region, which is defined as the area where the maximum phase error is less than $\pi/8$, i.e., $\bm{D}_{FF}=\{\bm{p}|\Delta\varphi(\bm{p})<\pi/8\}$. When the $l$-th scatter is located in the FF region, the path between the $l$-th scatter and the $n$-th RIS element can be modeled by
@@ II-C Localization Protocol
We propose an RIS-enabled hybrid NF and FF source localization protocol, which improves localization accuracy by adaptively localizing the users and optimizing the RIS phase shifts based on the estimated locations. The localization process is divided into $C$ cycles with cycle duration being $\delta$. Each cycle contains three steps: transmission, localization, and optimization. The process of the localization protocol is illustrated in Fig. 2.
@@ II-C1 Transmission
In this step, the users send signals to the RIS for $\delta_{T}$ seconds, which are reflected to the BS. Let $\bm{g}^{(c)}=[g^{(c)}_{1},...,g^{(c)}_{K}]^{T}\in\mathbb{C}^{K\times 1}$ denote the signal received by the BS in the transmission step of the $c$-th cycle. The RIS phase shifts are set randomly in the first cycle, while in the following $C-1$ cycles, the phase shifts are selected based on the optimization results in the previous cycle, which will be described in the optimization step.
@@ II-C2 Localization
In the next $\delta_{E}$ seconds, the BS estimates users’ locations using the received signals. Specifically, in the $c$-th cycle, based on the received signals $\bm{G}^{(c)}=[\bm{g}^{(1)},...,\bm{g}^{(c)}]^{T}\in\mathbb{C}^{K\times c}$, we determine whether the users and the scatters are located in the NF or the FF region and jointly estimate their locations $\bm{P}^{(c)}=[\bm{p}_{1}^{(c)},...,\bm{p}_{K}^{(c)}]$ and $\bm{Q}^{(c)}=[\bm{q}_{1}^{(c)},...,\bm{q}_{L}^{(c)}]$. This step focuses on the problem of accurately locating users given the received signals and the RIS phase shifts. The details of the localization algorithm are introduced in Sec. III.
@@ II-C3 Optimization
In summary, two parts of the high-level problem need to be solved: localization and optimization, which iterate with each other to achieve the optimal RIS phase shifts and higher localization accuracy. Two algorithms are designed to solve these problems respectively.
@@ III-A Localization Problem Formulation
where $\bm{g}^{(c)}_{k}=[g_{k}^{(1)},...,g_{k}^{(c)}]^{T}\in\mathbb{C}^{c\times 1}$ is the $k$-th row of received signal $\bm{G}^{(c)}$, i.e. the received signals of the $k$-th user in the previous $c$ cycles. Here $\bm{B}^{(c)}=[\bm{\beta}^{(1)},...,\bm{\beta}^{(c)}]\in\mathbb{C}^{N\times c}$ is the phase shifts of the RIS in the previous $c$ cycles, $\bm{P}=[\bm{p}_{1},...,\bm{p}_{K}]$ is the locations of the users, $\overline{\bm{Q}}$ is the ground truth of scatter locations, and $\|\cdot\|_{2}$ denotes the $l_{2}$-norm. It can be observed from (10a) that the objective function is affected by the accuracy of user location estimation. By solving (P1), we can minimize the estimation error of the user location and eliminate the interference caused by the scattering paths.
@@
$\bm{Z}_{near}=[[R_{1},\theta_{1},\phi_{1}]^{T},...,[R_{S},\theta_{S},\phi_{S}]^{T}].$ (11)
@@
$\bm{F}_{near}=[\bm{b}(R_{1},\theta_{1},\phi_{1}),...,\bm{b}(R_{S},\theta_{S},\phi_{S})]\in\mathbb{C}^{N\times S},$ (12)
where $\bm{b}(R_{i},\theta_{i},\phi_{i})\in\mathbb{C}^{N\times 1}$ is the steering vector given the location $[R_{i},\theta_{i},\phi_{i}]^{T}\in\bm{D}_{NF}$.
FF case: In the FF region, only the angles $\theta,\phi$ are sampled. The range $R$ is not sampled because it does not affect the FF steering vector and it cannot be estimated in a similar method to NF case. The primary reason for this difference lies in the angular relationships between the user and the RIS elements. Specifically, in the NF, the angles between the user and each RIS element are different, leading to different phase shift changes for paths from the user to each RIS element when the range varies. These differences in phase shift changes result in varying received signals, which allows for the estimation of the range $R$. In contrast, for FF users, since the angles between the user and each RIS element are almost the same, when the range changes, the phase shift changes for the path from the user to each RIS element are the same. This uniform phase shift leads to an overall phase change in the received signal, hence range estimation cannot be achieved.
@@
where $\bm{a}(\theta_{i},\phi_{i})$ is the steering vector of the FF location $[\theta_{i},\phi_{i}]$.
@@
Note that $\bm{u}_{k}$ can be directly solved base on traditional sparse recovery method like OMP. However, the user location estimated in this way is not accurate enough due to interference of scattering paths. To mitigate this interference and improve user localization accuracy, (P1) is approximated as
@@
$s.t.$ $\|\bm{u}^{\bm{P}}_{k}\|_{0}=1,\forall k,$ (16b)
where $\|\cdot\|_{0}$ denotes the $l_{0}$-norm, and $\overline{\bm{u}^{\bm{Q}}_{k}}$ is the ground truth of ${\bm{u}^{\bm{Q}}_{k}}$. Here, $\bm{\Lambda}^{k}=(\bm{B}^{(c)})^{T}\text{diag}\{\bm{h}^{A}_{k}\}\bm{F}s_{k}\in\mathbb{C}^{c\times M}$ is the atom signals, which are the linear transformation of the atom channels $\bm{F}$ by considering the effect of RIS phase shift $\bm{B}^{(c)}$, the RIS-BS channel $\bm{h}^{A}_{k}$, and the transmit signal $s_{k}$.
@@ III-B Localization Algorithm
From (P1’), it is evident that we cannot directly estimate the user’s location since the ground truth of the scatter locations is unknown. Therefore, we need to estimate the locations of the scatters. However, considering that scattering paths are typically weaker than direct user-RIS path, it is difficult to estimate the scatter locations directly from the received signals. To tackle this issue, we leverage the idea of successive interference cancellation (SIC) that first estimates the stronger signal components, i.e., the direct user-RIS paths, and then estimates the weaker scattering paths.
@@
12:     if $E^{r}<\gamma E$ then
13:        break
14:     end if
15:  end for
16:  for $k=1,...,K$ do
17:     Calculate the new residual signal by (28)
18:     Obtain the estimated location of $k$-th user by (30)
19:  end for
20:  Return: $\bm{P}^{(c)}=[\bm{p}_{1}^{(c)},...,\bm{p}_{K}^{(c)}]$

### v2

@@ I-A Literature Review and Motivation
Traditional HMIMO-based localization works mainly fall into two categories: far-field (FF)[5, 6, 7, 8, 9, 10, 11] and near-field (NF) localization schemes [12, 13, 14, 15, 16, 17]. In the FF region, the distances between the HMIMO and users are sufficiently large, and the signal channel is effectively described using a plane wave model. In [5], the fundamental bounds and localization feasibility condition are examined for FF localization in orthogonal frequency division multiplexing (OFDM) systems aided by HMIMO. The authors in [6] considered a single-input-single-output (SISO) HMIMO-aided localization system with imperfect HMIMO geometry information, and derived the misspecified Cramér-Rao bound (CRB). In [7], the positioning estimation error and orientation estimation error are investigated for a HMIMO-aided multiple-input multiple output system. In [8], the authors analyzed the CRB for a localization scenario that both the line-of-sight (LOS) path and the reflected path via the HMIMO exist simultaneously. Two HMIMO phase design methods were proposed and a maximum likelihood (ML) direction of arrival (DOA) estimator was designed in [8]. In [9], the authors considered FF user localization aided by millimeter wave (mmWave) HMIMO, and a ML location estimator and two sub-optimal estimators with lower complexity were derived. In [10], the localization method leveraged the dynamic regulation of HMIMO to cancel non-line-of-sight signals at the receiver and preserve LOS signals for precise indoor localization. HMIMO-based localization without base stations (BSs) was also investigated in [11], where the user transmitted the signal and received the signal reflected from the HMIMO.
@@
HMIMO-based NF localization schemes have attracted much attention recently due to the increasing interest in large-scale HMIMO. Specifically, to improve beam control capability, the scale of HMIMO is enlarged, leading to the expansion of the NF region[18]. Different from the FF case, the spherical wave model has to be employed to characterize the NF channel because the signal transmission characteristics of the NF are determined by both the range and angle of the user location. The authors in [12] and [13] considered a SISO localization scenario for NF user, and the CRB is derived. In [14], a HMIMO functioning as a lens receiver was utilized to achieve NF localization. The authors in [15] developed a received signal strength-based localization method for NF users and a RIS phase shift adjustment method to enhance accuracy. In [16], the authors utilized light detection and ranging to estimate scatter locations and improve NF user localization accuracy. A signaling and positioning method is proposed in [17], and the coverage capability under different levels of obstruction of the HMIMO is analyzed.
@@
In practical scenarios, users may exist in both the NF and FF regions of the HMIMO. Existing algorithms for pure NF or FF users are therefore not always applicable. Specifically, FF localization methods suffer from low localization accuracy for NF users due to the significant error in describing NF channels using the plane wave model. Meanwhile, NF localization techniques are not ideal for FF users because the coupling of range and angle in the spherical model brings extra complexity, which has two main effects: first, it prolongs the running time of the algorithm; second, the more complex NF model hinders convergence of the method and leads to accuracy degradation. In addition, the localization gain for the FF users brought by the NF model is limited. Hence, a localization scheme is required to adapt to both the NF and FF users.
I-B Main Contributions
@@
Several challenges in this scenario need to be addressed. First, it is unknown whether each user or scatter is in the NF or the FF region, leading to difficulty in designing the localization algorithm. Second, the channels of different users are coupled due to shared RIS phase shifts and common scatters. To tackle the above challenges efficiently, a hybrid NF and FF localization algorithm is proposed, in which we first sample the NF and FF regions separately, and then compare the received signals with the reconstructed signals at the sampled locations to locate the users and scatters. The locations of the users and scatters are jointly estimated by considering the channel coupling effect. Then, a RIS phase shift optimization problem is formulated to minimize the sum of the Cramér-Rao bounds (CRBs) of the user localization errors. A complex circle manifold-based method is proposed to solve the optimization problem with the coupled CRBs.
@@
I-C Organization
The rest of this paper is organized as follows. The localization scenario, the signal model, and the localization protocol are provided in Sec. II. In Sec. III, we propose the localization algorithm. Then, the RIS phase shift optimization algorithm is designed in Sec. IV. The performance of the proposed method is analyzed in Sec. V. In Sec. II, we present the simulation results and discussions. Finally, conclusions are drawn in Sec. VII.
@@ II System Model
TABLE I: Major Notations
Notation Description
$N$ Number of RIS elements
$K$ Number of users
$L$ Number of scatters
$C$ Number of localization cycles
$\bm{\beta}$ RIS phase shifts
$\bm{B}^{(c)}$ RIS phase shifts for $c$ cycles
$\bm{h}_{k}$ Channel between the $k$-th user and the BS
$\bm{h}^{A}_{k}$ Channel between the RIS and the BS
$\bm{h}^{t}_{k}$ Channel between the RIS and the $k$-th user
$s_{k}$ The transmitted signal by the $k$-th user
$\bm{p}_{k}$ Location of the $k$-th user
$\bm{Q}$ Locations of all scatters
$\bm{g}^{(c)}_{k}$ The received signal of the $c$-th cycle
$\bm{Z}$ The sampled locations
$\bm{F}$ Atom channel in the NF and FF regions
$\bm{u}_{k}$ Gains of the atom channels for the $k$-th
$\bm{u}_{k}^{\bm{P}}$ The direct path component in $\bm{u}_{k}$
$\bm{u}_{k}^{\bm{Q}}$ The scattering path component in $\bm{u}_{k}$
$\bm{\Lambda}^{k}$ Atom signals for the $k$-th user
$\bm{D}_{NF}$, $\bm{D}_{FF}$ NF and FF region
$D_{R}$, $D_{\theta}$, $D_{\phi}$ CRBs for range and angle
@@ II-A Localization Scenario
As shown in Fig. 1, we consider a localization scenario consisting of $K$ users, a RIS, and a BS [^1] In this work, we use reflective reconfigurable intelligent surface (RIS) as an example to implement HMIMO, and it can be easily replaced with other surfaces, such as transmissive RIS and reconfigurable holographic surface.. The users are stationary or low-speed compared with the localization time. Users and the BS are each equipped with a single antenna. The RIS containing $N=N_{1}\times N_{2}$ elements is placed on the Y-Z plane, where the center of the RIS is at the origin of the coordinate system. There also exist obstacles, which reflect the signals and create multipath channels. We assume the line-of-sight path between each user and the BS is blocked, and there only exist the reflection paths via the RIS.
During the localization process, each user sends a narrowband signal to the RIS, and the RIS reflects it to the BS. To avoid multi-user interference, frequency division multiplexing (FDM) is for different users, i.e. the $k$-th user sends the signal over frequency $f_{k}$, where $f_{k}\neq f_{k}^{\prime}$ ($k\neq k^{\prime}$). The BS then estimates the locations of the users using the received signals. We assume the locations of the RIS and the BS are known, while the locations of the users and scattering centers in the obstacles [19] are unknown. Whether the users are located in the NF or the FF regions of the RIS is also unknown.
@@ II-B1 Channel Models for the NF Region
$\!\!\!\bm{b}(R_{k},\theta_{k},\phi_{k})\!=\!\!\left[\exp\!\left(\!-j\dfrac{2\pi}{\lambda}d_{k1}^{t}\!\right)\!,\dots,\exp\!\left(\!-j\dfrac{2\pi}{\lambda}d_{kN}^{t}\!\right)\!\right]^{T}\!\!\!\!.\!\!$ (4)
@@
$\!\Delta\varphi(\bm{p})\!=\!\mathop{\max}\limits_{n}\frac{2\pi}{\lambda}\!\left(d_{n}^{t}-(R-y_{n}\sin\theta\sin\phi-\!z_{n}\cos\theta)\right),$ (5)
@@
Similarly, we model the paths via the scatters. Specifically, when the $l$-th scatter is located in the NF region, the path between the $l$-th scatter and the RIS is given by
@@ II-B2 Channel Models for the FF Region
where $\bm{a}(\theta_{k},\phi_{k})$ is the FF steering vector and is given as [23] [24]
@@
In (7), $\bm{D}_{FF}$ represents the FF region, which is defined as the area where the maximum phase error is less than $\pi/8$, i.e., $\bm{D}_{FF}=\{\bm{p}|\Delta\varphi(\bm{p})<\pi/8\}$. When the $l$-th scatter is located in the FF region, the path between the $l$-th scatter and the RIS can be modeled by
@@ II-C Localization Protocol
We propose an RIS-enabled hybrid NF and FF localization protocol, which improves localization accuracy by iteratively localizing the users and optimizing the RIS phase shifts based on the estimated locations. The localization process is divided into $C$ cycles with cycle duration being $\delta$. Note that cycle duration $\delta$ can be adjusted for different systems. Each cycle contains three steps: transmission, localization, and optimization. The process of the localization protocol is illustrated in Fig. 2.
@@ II-C1 Transmission
In this step, the users send signals to the RIS, which are reflected to the BS. Let $\bm{g}^{(c)}=[g^{(c)}_{1},...,g^{(c)}_{K}]^{T}\in\mathbb{C}^{K\times 1}$ denote the signal received by the BS in the $c$-th cycle. The RIS phase shifts are set randomly in the first cycle, while in the following $C-1$ cycles, the phase shifts are selected based on the optimization results in the previous cycle, which will be described in the optimization step.
@@ II-C2 Localization
In the next step, the BS estimates users’ locations using the received signals. Specifically, in the $c$-th cycle, based on the received signals $\bm{G}^{(c)}=[\bm{g}^{(1)},...,\bm{g}^{(c)}]^{T}\in\mathbb{C}^{K\times c}$, we determine whether the users and the scatters are located in the NF or the FF region and jointly estimate their locations $\bm{P}^{(c)}=[\bm{p}_{1}^{(c)},...,\bm{p}_{K}^{(c)}]$ and $\bm{Q}^{(c)}=[\bm{q}_{1}^{(c)},...,\bm{q}_{L}^{(c)}]$. This step focuses on the problem of accurately locating users given the received signals and the RIS phase shifts. The details of the localization algorithm are introduced in Sec. III.
@@ III-A Localization Problem Formulation
where $\bm{g}^{(c)}_{k}=[g_{k}^{(1)},...,g_{k}^{(c)}]^{T}\in\mathbb{C}^{c\times 1}$ is the received signals of the $k$-th user in the previous $c$ cycles, i.e., the $k$-th row of $\bm{G}^{(c)}$. Here $\bm{B}^{(c)}=[\bm{\beta}^{(1)},...,\bm{\beta}^{(c)}]\in\mathbb{C}^{N\times c}$ is the phase shifts of the RIS in the previous $c$ cycles, $\bm{P}=[\bm{p}_{1},...,\bm{p}_{K}]$ is the locations of the users, $\overline{\bm{Q}}$ is the ground truth of scatter locations. It can be observed from (10a) that the objective function is affected by the accuracy of user location estimation. By solving (P1), we can minimize the estimation error of the user location and eliminate the interference caused by the scattering paths.
@@
$\bm{Z}_{near}=\left[\left[R_{1},\theta_{1},\phi_{1}\right]^{T},...,\left[R_{S},\theta_{S},\phi_{S}\right]^{T}\right].$ (11)
@@
$\bm{F}_{near}=\left[\bm{b}(R_{1},\theta_{1},\phi_{1}),...,\bm{b}(R_{S},\theta_{S},\phi_{S})\right]\in\mathbb{C}^{N\times S},$ (12)
where $\bm{b}(R_{i},\theta_{i},\phi_{i})$ is the NF steering vector given the location $[R_{i},\theta_{i},\phi_{i}]^{T}$.
FF case: In the FF region, only the angles $\theta,\phi$ are sampled. The range $R$ is not sampled because it does not affect the FF steering vector and it cannot be estimated similarly to NF case. The primary reason lies in the angular relationships between the user and the RIS elements. Specifically, in the NF, the angles between the user and each RIS element are different. This leads to different phase shift changes for paths from the user to each RIS element and varied received signals when the range varies, which allows for the estimation of the range $R$. In contrast, for FF users, since the angles between the user and each RIS element are almost the same, when the range changes, the phase shift changes for the path from the user to each RIS element are the same. This uniform phase shift leads to an overall phase change in the received signal, hence range estimation cannot be achieved.
@@
where $\bm{a}(\theta_{i},\phi_{i})$ is the FF steering vector of the angle $[\theta_{i},\phi_{i}]$.
@@
Since the number of users and scatters are generally much smaller than the number of atom channels, $\bm{u}_{k}$ can be directly solved base on traditional sparse recovery method like orthogonal matching pursuit (OMP). However, the user location estimated in this way is not accurate enough due to interference of scattering paths. To mitigate this interference and improve user localization accuracy, (P1) is approximated as
@@
$s.t.$ $\left\|\bm{u}^{\bm{P}}_{k}\right\|_{0}=1,\forall k,$ (16b)
where $\overline{\bm{u}^{\bm{Q}}_{k}}$ is the ground truth of ${\bm{u}^{\bm{Q}}_{k}}$. Here, $\bm{\Lambda}^{k}=(\bm{B}^{(c)})^{T}\text{diag}\{\bm{h}^{A}_{k}\}\bm{F}s_{k}\in\mathbb{C}^{c\times M}$ is the atom signals, which are the linear transformation of the atom channels $\bm{F}$ by considering the effect of RIS phase shift $\bm{B}^{(c)}$, the RIS-BS channel $\bm{h}^{A}_{k}$, and the transmit signal $s_{k}$.
@@ III-B Localization Algorithm
From (P1’), it is evident that we cannot directly estimate the user’s location since the scatter locations are unknown. Therefore, we need to estimate the locations of the scatters. However, considering that scattering paths are typically weaker than direct user-RIS path, it is difficult to estimate the scatter locations directly from the received signals. To tackle this issue, we leverage the idea of successive interference cancellation (SIC) that first estimates the stronger direct user-RIS paths, and then estimates the weaker scattering paths.
@@
12:     if $E^{r}<\gamma E$ then break end if
13:  end for
14:  for $k=1,...,K$ do
15:     Calculate the new residual signal by (28)
16:     Obtain the estimated location of $k$-th user by (30)
17:  end for
18:  Return: $\bm{P}^{(c)}=[\bm{p}_{1}^{(c)},...,\bm{p}_{K}^{(c)}]$